OpenStudy (anonymous):

You plan to drive north from city A to to town B and then continue on to city C. The distance between city A and town B is 39 miles, and the distance between town B and city C is 99 miles * - assuming you following a straight driving path, after how many miles of driving will you reach the midpoint between city A and City C ? * if you drive an average of 46 mi/hr, how long will ot take to drive from city A to city C?

OpenStudy (dumbcow):

midpoint = (99+39)/2 = 69 distance = rate*time 138 = 46t t = 138/46 = 3